Tickets for the homecoming dance cost $20 for a single ticket or $35 for a couple. Ticket sales totaled $2280, and 128 people attended. How many tickets of each type were sold?
Answer by stanbon(75887)

Tickets for the homecoming dance cost $20 for a single ticket or $35 for a couple. Ticket sales totaled $2280, and 128 people attended. How many tickets of each type were sold?
-------
Quantity Eq.::: s + 2d = 128
Value Eq::::: 20s +35d = 2280
--------------------------------
20s + 40d = 20*128
20s + 35d = 2280
--------------------
Subtract and solve for "d":
5d = 280
d = 56 (# of double tickets sold)
-----
Solve for "s":
s + 2*56 = 128
s + 112 = 128
s = 16 (# of single tickets sold)
===================================
Cheers,
Stan H.
============